OpenStudy (anonymous):

In the function f(x) = 4(x2 − 6x + ____) + 20, what number belongs in the blank to complete the square?

OpenStudy (mathstudent55):

Take half of the x-term coefficient and square it.

OpenStudy (mathstudent55):

The x-term coefficient is -6. Take half of -6, and square it. What do you get?

OpenStudy (anonymous):

Half of it so like -3?

OpenStudy (mathstudent55):

Yes. Now square it.

OpenStudy (anonymous):

ok its -1.73

OpenStudy (anonymous):

or well - 1.7

OpenStudy (mathstudent55):

Not square root. Square it. \((-3)^2 = (-3)(-3) = 9\)

OpenStudy (anonymous):

Oh square it XD

OpenStudy (mathstudent55):

\(\sqrt 3 \approx 1.732\) <----- not this \((-3)^2 = 9\) <---- this is what you do

OpenStudy (anonymous):

Ok understood!

OpenStudy (mathstudent55):

The answer is 9. Just so you know, this method (take half of the x-term coefficient and square it) works for completing the square, but it only works when the coefficient of the x^2 term is 1. In your case, you do have x^2 - 6x, so you see that the coefficient of the x^2 term is 1.

OpenStudy (anonymous):

Thanks

OpenStudy (mathstudent55):

You're welcome.